# Mapping atmospheric features of the planetary-mass brown dwarf SIMP 0136 with JWST NIRISS

### v1

@@ 6 Mapping
We may infer some brightness geometry of stars and planes via observed light curve variations. This inference is complicated by the fact that the brightness-to-map problem is degenerate: one brightness map results in a single light curve, but one light curve can be produced by potentially infinite map variants. Knowledge of rotational period, inclination and limb darkening can help to constrain some parameters.
@@
In this work we use two complementary techniques to map SIMP 0136: sinusoidal (Cowan & Agol, 2008; Cowan et al., 2013), and spherical harmonic mapping (Cowan et al., 2013; Luger et al., 2019). The first method is algebraically simple and assumes a north-south symmetry of the object. The second uses spherical harmonic basis maps.
@@
If the thermal map is North-South symmetric, then odd harmonics do not have a phase function signature. Instead, our observations of SIMP 0136 exhibit significant power at the 3rd mode (see Figure 6b), indicating North-South asymmetry and inclined rotation (Cowan et al., 2013).
@@
Spherical harmonics within the Starry framework (Luger et al., 2019) allow to account for the inclination of the object. We provide a multivariate Gaussian prior on the spherical harmonics coefficients with a mean and covariance.
@@
The results from fitting the m band lightcurve with various mapping methods are shown in Fig. 6a. A sinusoidal fit with 4 modes, excluding the 3rd mode, is represented by the dashed green line. The maximum number of modes was selected using Bayesian Information Criterion (BIC). This approach failed to capture the distinct shape of the phase curve, nor similar features in other bands. If we include the odd harmonics (dashed blue line), we get a reasonable fit. This means that SIMP 0136 has a North-South asymmetric map.
@@
Spherical harmonic decomposition accounts for inclination and produces 2D maps. Using BIC we constrained number of -degrees to 4, that is 25 Ylm harmonics in total. We provided 2 sets of priors on coefficients: flat with uniform variance, and a spot-band structure. The resulting maps can be seen in Fig. 6c.1 and c.2. The asymmetric checker-board pattern is the most favourable for flat priors. A more plausible map is obtained with a spot-band prior. The fact that these maps produce identical lightcurves is a testament to the extreme mapping degeneracy (Cowan et al., 2013).
7 Connecting vertical structure to lightcurve morphology
@@
SIMP 0136’s spectrally inhomogeneous variability indicates multiple layers with distinct atmospheric processes and dynamics. The enhanced NIR spectrophotometric variability observed at the L/T spectral transition has been demonstrated to be best fit by Fourier modes for objects such as WISE 1049B (aka Luhman 16B), VHS 1256 b, and SIMP 0136 (e.g., Apai et al., 2017, 2021; Fuda et al., 2024; Zhou et al., 2020a; Plummer et al., 2024). In this section, we employ a contribution function (and our previous harmonic analysis above) to map spectral wavelength ranges to pressure levels and create 2D (pressure vs time) flux variability maps which can be decomposed to determine variability sources at each atmospheric layer.
@@
First, we fit the multi-spectral lightcurves with Fourier models using Imber (Plummer & Wang, 2022, 2023) which employs dynamic nested sampling (Skilling, 2004, 2006) via the open source Python code Dynesty (Speagle, 2020). Considering the spectral band from , we use bins to plot 36 light curves (see Figure 1 for an example with 0.2 bins). Qualitatively, three characteristic light curve shapes can be seen in these results, broadly matching the findings of Biller et al. (2024) and Chen et al. (2025) for WISE 1049B and McCarthy et al. (2025) for SIMP 0136 which found 3, 4 and 9 characteristic light curve shapes, respectively, with JWST/NIRSpec data. For each binned light curve, we select the best-fit Fourier model based on the Bayesian Information Criterion.
@@
To create vertical maps, we use the best-fit Fourier models and a 1150 K, (Vos et al., 2023; Gagné et al., 2017), cloud-free Sonora Bobcat substellar atmospheric model (Marley et al., 2021) and Picaso (Batalha et al., 2019). At each time-step, we scale 0.05 m interval spectral bins in the contribution function by the corresponding normalized light curves from Figure 1. We then sum the total emission flux at each pressure level across all wavelengths. This creates a 1D vector of fluxes as a function of pressure. We repeat this process at every time step and then normalize the flux at each pressure level. The resulting vertical flux variability maps (see Figure 7) suggest three distinct layers, which match the qualitative light-curve morphology groups seen in Figure 1.
@@
The deepest layer () corresponds to (Figure 7a), includes higher-order harmonics (, where is the wave number), and corresponds to the spectral region where the models predict cloud variability (e.g., McCarthy et al., 2025). We interpret higher-order harmonics at this pressure level as forsterite (Mg2SiO4) cloud modulation, in line with Vos et al. (2023), McCarthy et al. (2024), and Plummer et al. (2024).
The transition layer () at 1 bar corresponds to the expected pressure level of Mg2SiO4 cloud tops (e.g., Vos et al., 2023; McCarthy et al., 2024) and exhibits a transition from waves including complex, higher-order harmonic waves to those dominated by lower-order harmonics. The layer contains bright spots correlated with the suspected cloud features at the top of the deeper layer. This layer corresponds to a prominent water absorption band and we suspect the bright spot in Figure 7b depicts either localized heating or variations in observed water absorption tied to the presence of forsterite clouds, either due to muting of spectral features or the depletion of water as part of the thermochemical reaction that produces forsterite clouds (Visscher et al., 2010).
The highest altitude layer () with pressures mbar (see Figure 7c) typically contains lower-order harmonics (). This high layer’s wavelength range includes both water and carbon monoxide (CO) absorption bands. As can be seen by comparison to Figure 7c, the upper layer’s flux variability is in phase with the bright spot in the transition layer at bar.
8 Regional spectra retrieval
@@
In our case, the extracted regional spectra and their corresponding atmospheric retrievals are unreliable due to degeneracies inherent in 2D mapping. We also encountered difficulties due the absence of grid-based flux in the forward model. As already shown by Cubillos et al. (2021), the grid flux normalization is not trivial. Nevertheless, some results can be interpreted. The temperature structure varies significantly across different regions, suggesting that variability is primarily due to temperature inhomogeneity. Some species are well-constrained and consistent between regions: water, methane, alkalies, iron hydride. However, regional spectra show higher water, alkali, and methane abundances than time-averaged spectrum. At the same time, CO and carbon oxide, ammonia, chromium hydride have very different and uncertain posterior distributions. This could suggest that carbon chemistry is not the main source of variability. Additional details on the regional spectral retrieval attempts are provided in Appendix B.
@@
9 Doppler Tomography
While the spectral resolution of NIRISS/SOSS is modest compared to ground-based precision radial velocity (pRV) spectrographs, the instrument’s stability enables meaningful Doppler measurements. The SOSSISSE framework describes the data as fractional variations relative to a mean flux, presented as a residual map. To measure velocity shifts, we adopt the approach of Bouchy et al. (2001), in which radial velocity shifts are determined by projecting the residuals onto the derivative of a high-SNR template. Although originally developed for the pRV field, this method is applicable to any spectroscopic time series. The Bouchy et al. (2001) formalism also provides associated uncertainties (see their Eq. 13).
Given the low spectral resolution of SOSS and its limited constraints on radial velocities, we do not attempt to divide the domain into wavelength bins as is done with flux variability. Instead, we consider the integral of Eq. 9 from Bouchy et al. (2001) across the 1.00–2.35 m range. The redder portion of the spectrum is excluded due to its markedly different light curve behavior.
Figure 8 shows the resulting radial velocity time series, revealing correlated variations at the km/s level, characterized by a three-harmonic signal. Per-point uncertainties are approximately 1 km s-1. The signal is in phase with the time derivative of the flux, consistent with the framework described by Aigrain et al. (2012), in which RV variability for a rotating object (neglecting convective inhibition) scales with the product of the flux and its time derivative—the so-called term in activity-related RV corrections.
@@ 9 Doppler Tomography
These results underscore an unexplored opportunity in JWST time-series spectroscopy of brown dwarfs. The high precision of molecular-band flux variability measurements—informing cloud structures at different altitudes—can be complemented by RV-based constraints on longitudinal and latitudinal features. While phase-resolved light curves constrain longitudinal structures, RV amplitudes are particularly sensitive to low-latitude features, providing strong priors for brightness map reconstructions. Although the SOSS data for SIMP 0136 lacks sufficient SNR to allow subdivision by bandpass, higher-resolution modes such as NIRSpec’s G140H or G235H are of particular interest. For a profile limited by instrumental resolution, RV uncertainty scales with the 3/2 power of resolution at a fixed flux level. Thus, moving from (SOSS) to (G140H/G235H) yields a theoretical sevenfold improvement in SNR, enabling subdivision into several tens of spectral segments while retaining km/s-level RV precision.
10 discussion and conclusion
In this work, we presented a multifaceted analysis of time-resolved spectroscopy of the planetary mass object SIMP J01365662+093347. The data were acquired with JWST NIRISS/SOSS and reduced with SOSSISSE. As expected from JWST’s stability, and amply demonstrated with transit work, the data is far beyond what could be obtained through ground-based observations of BDs. This level of detail opens a new window into BD atmospheric dynamics.
PCA showed that there are at least 3 different spectral regions rotating in and out of view. Indeed, not a single atmospheric model could fit observed spectrum, and a mixture of 6 Diamondback models is needed to adequately fit the time-averaged spectrum. Projected interpolated Diamondback models on the PCP suggest that disk-integrated physical parameters vary in time. These parameters are temperature, metallicity, and sedimentation efficiency. The fact that we have small spectral variations and change in integrated physical parameters in time ( K, f, ), suggests that SIMP 0136 has small-scale cloud patches of different temperature, as predicted by brown dwarf GCM results (Tan & Showman, 2021b). Discrepancy in the reported temperatures of Diamondback models does not affect the magnitude of temperature variability as it only creates a linear bias. The inhomogeneity in the integrated effective metallicity may arise from different amounts of chemical disequilibrium in each spectral region, e.g., differences in the CO/CH4 ratios among cloud layers.
@@
Odd harmonics in the lightcurve of SIMP 0136 and the results from spherical harmonics mapping imply North-South asymmetry, combined with the previously noted inclined rotation. The 2D maps inferred from a single rotation remain degenerate. Nevertheless, lightcurve morphology in combination with Sonora Bobcat contribution functions reveal multidimensional structure of observed variability. We find that the spatial distribution of forsterite clouds below the 1-bar level is anti-correlated with the abundances of water and CO at higher altitudes. This is evidenced by a decrease in disk-integrated emissivity from forsterite layers, accompanied by an increase in emissivity from layers associated with water and CO absorption bands. North-South asymmetry of the brightness map in the domain could directly arise from the spatial distribution of small-scale forsterite cloud patches. It is not guaranteed that this weather pattern is a permanent feature on SIMP 0136: multi-period back to back observations could show how this morphology evolves.
@@
The fact that the meteorology of SIMP 0136 evolves is confirmed by several ground- and space-based observations at different epochs. In Artigau et al. (2009), -band (m) photometric observations at Observatoire du Mont-Mégantic in 2008 reveal how the power of Fourier modes varies across nights. Phase-folded lightcurves from September 18th and 19th can be well explained by the first 2 modes, while phase curves of September 16th and 21st require an odd 3rd mode harmonic. Further -band observations at Lowell Observatory in 2016 (Croll et al., 2016) show consistent single-peak lightcurves across 4 rotations on the night of November 10th. On the night of November 12th, SIMP 0136 shows almost no variability in -band, indicating that meteorological features have longitudinally homogenized. The next night, however, sawtooth-like variations are consistently observed again, but in an anti-correlation with the night of November 10th, which suggests a slightly denser distribution of clouds in one hemisphere. Plummer et al. (2024) analyzed CFHT -band observations obtained in 2012: the lightcurve from October 14th is described by 3 Fourier modes; however, the following nights observations alternated between Y, J, H, and K bands, limiting the cadence and only allowing 2 modes to be inferred for each band. A phase shift between H-K and the J-H and H-K color indices was found, which might have indicated different cloud layers. Lastly, we compare our observations with JWST GO #3548 NIRSpec/BOTS time-series spectroscopy analyzed by McCarthy et al. (2025). The binned lightcurve in 0.9 – 2.1 m range is well explained by 3 Fourier modes, with the 3rd mode having equal power as the 2rd mode, just like in our observations. Yet, 37 hours after our observations, the 1st mode has evolved and increased its contribution, leading to a different lightcurve morphology between the 2 visits, which is not explained by a simple shift in phase.
@@
We explored a method of extracting and retrieving spectra using spherical harmonics mapping at different wavelengths. This method is valuable in inferring chemical abundances and thermal profiles of potential hot and cold spots of the brightness maps. We performed spectral retrievals on 6 regions of interest but the results are not consistent. This is due to the degeneracy of 2D mapping. Observations spanning multiple rotations or Doppler tomography could decrease these degeneracies.
@@
We also use JWST medium resolution spectroscopy for Doppler tomography of SIMP 0136. These measurement could eventually help constrain brightness maps and improve mapping fidelity. The search for a discrepancy between Doppler signal and lightcurve Fourier modes would benefit from higher spectral resolution observations.

### v2

@@ VI Mapping
We may infer some brightness geometry of stars and planets via observed lightcurve variations. This inference is complicated by the fact that the brightness-to-map problem is degenerate: one brightness map results in a single lightcurve, but one lightcurve can be produced by potentially infinite map variants. Knowledge of rotational period, inclination, and limb darkening can help to constrain some parameters.
@@
In this work we use two complementary techniques to map SIMP 0136: sinusoidal (Cowan & Agol, 2008; Cowan et al., 2013) and spherical harmonic mapping (Cowan et al., 2013; Luger et al., 2019). The first method is algebraically simple and assumes a north–south symmetry of the object. The second uses spherical harmonic basis maps.
@@
If the thermal map is north–south symmetric, then odd harmonics do not have a phase function signature. Instead, our observations of SIMP 0136 exhibit significant power at the third mode (see Figure 6b), indicating north–south asymmetry and inclined rotation (Cowan et al., 2013).
@@
Spherical harmonics within the Starry framework (Luger et al., 2019) allow us to account for the inclination of the object. We provide a multivariate Gaussian prior on the spherical harmonics coefficients with a mean and covariance.
@@
The results from fitting the –m band lightcurve with various mapping methods are shown in Figure 6a. A sinusoidal fit with four modes, excluding the third mode, is represented by the dashed green line. The maximum number of modes was selected using the Bayesian Information Criterion (BIC). This approach failed to capture the distinct shape of the phase curve, nor similar features in other bands. If we include the odd harmonics (dashed blue line), we get a reasonable fit. This means that SIMP 0136 has a north–south asymmetric map.
@@
Spherical harmonic decomposition accounts for inclination and produces 2D maps. Using BIC, we constrained the number of -degrees to four, that is, 25 Ylm harmonics in total. We provided two sets of priors on coefficients: flat with uniform variance, and a spot-band structure. The resulting maps can be seen in Figure 6c.1 and c.2. The asymmetric checkerboard pattern is the most favorable for flat priors. A more plausible map is obtained with a spot-band prior. The fact that these maps produce identical lightcurves is a testament to the extreme mapping degeneracy (Cowan et al., 2013).
VII Connecting vertical structure to lightcurve morphology
@@
SIMP 0136’s spectrally inhomogeneous variability indicates multiple layers with distinct atmospheric processes and dynamics. The enhanced near-infrared (NIR) spectrophotometric variability observed at the L/T spectral transition has been demonstrated to be best fit by Fourier modes for objects such as WISE 1049B (aka Luhman 16B), VHS 1256 b, and SIMP 0136 (e.g., Apai et al., 2017, 2021; Fuda et al., 2024; Zhou et al., 2020a; Plummer et al., 2024). In this section, we employ a contribution function (and our previous harmonic analysis above) to map spectral wavelength ranges to pressure levels and create 2D (pressure-versus-time) flux variability maps, which can be decomposed to determine variability sources at each atmospheric layer.
@@
First, we fit the multi-spectral lightcurves with Fourier models using Imber (Plummer & Wang, 2022, 2023), which employs dynamic nested sampling (Skilling, 2004, 2006) via the open-source Python code Dynesty (Speagle, 2020). Considering the spectral band from , we use bins to plot 36 lightcurves (see Figure 1 for an example with 0.2 bins). Qualitatively, 3 characteristic lightcurve shapes can be seen in these results, broadly matching the findings of Biller et al. (2024) and Chen et al. (2025) for WISE 1049B and McCarthy et al. (2025) for SIMP 0136, which found 3, 4, and 9 characteristic lightcurve shapes, respectively, with JWST/NIRSpec data. For each binned lightcurve, we select the best-fit Fourier model based on the Bayesian Information Criterion.
@@
To create vertical maps, we use the best-fit Fourier models and a 1150 K, (Gagné et al., 2017; Vos et al., 2023), cloud-free Sonora Bobcat substellar atmospheric model (Marley et al., 2021) and Picaso (Batalha et al., 2019). At each time step, we scale 0.05 m interval spectral bins in the contribution function by the corresponding normalized lightcurves from Figure 1. We then sum the total emission flux at each pressure level across all wavelengths. This creates a 1D vector of fluxes as a function of pressure. We repeat this process at every time step and then normalize the flux at each pressure level. The resulting vertical flux variability maps (see Figure 7) suggest three distinct layers, which match the qualitative lightcurve morphology groups seen in Figure 1.
@@
The deepest layer (–) corresponds to (Figure 7(a)), includes higher-order harmonics (, where is the wavenumber), and corresponds to the spectral region where the models predict cloud variability (e.g., McCarthy et al., 2025). We interpret higher-order harmonics at this pressure level as forsterite (Mg2SiO4) cloud modulation, in line with Vos et al. (2023), McCarthy et al. (2024), and Plummer et al. (2024).
The transition layer (–) at corresponds to the expected pressure level of Mg2SiO4 cloud tops (e.g., Vos et al., 2023; McCarthy et al., 2024) and exhibits a transition from waves including complex, higher-order harmonic waves to those dominated by lower-order harmonics. The layer contains bright spots correlated with the suspected cloud features at the top of the deeper layer. This layer corresponds to a prominent water absorption band, and we suspect the bright spot in Figure 7(b) depicts either localized heating or variations in observed water absorption tied to the presence of forsterite clouds, either due to muting of spectral features or the depletion of water as part of the thermochemical reaction that produces forsterite clouds (Visscher et al., 2010).
The highest altitude layer (–) with pressures (see Figure 7(c)) typically contains lower-order harmonics (). This high layer’s wavelength range includes both water and carbon monoxide (CO) absorption bands. As can be seen by comparison to Figure 7(c), the upper layer’s flux variability is in phase with the bright spot in the transition layer at .
VIII Regional spectra retrieval
@@
In our case, the extracted regional spectra and their corresponding atmospheric retrievals are unreliable due to degeneracies inherent in 2D mapping. We also encountered difficulties due to the absence of grid-based flux in the forward model. As already shown by Cubillos et al. (2021), the grid flux normalization is not trivial. Nevertheless, some results can be interpreted. The temperature structure varies significantly across different regions, suggesting that variability is primarily due to temperature inhomogeneity. Some species are well constrained and consistent between regions: water, methane, alkalies, and iron hydride. However, regional spectra show higher water, alkali, and methane abundances than the time-averaged spectrum. At the same time, CO and carbon oxide, ammonia, and chromium hydride have very different and uncertain posterior distributions. This could suggest that carbon chemistry is not the main source of variability. Additional details on the regional spectral retrieval attempts are provided in Appendix B.
@@
IX Doppler Tomography
While the spectral resolution of NIRISS/SOSS is modest compared to ground-based precision radial velocity (pRV) spectrographs, the instrument’s stability enables meaningful Doppler measurements. The SOSSISSE framework describes the data as fractional variations relative to a mean flux, presented as a residual map. To measure velocity shifts, we adopt the approach of Bouchy et al. (2001), in which radial velocity shifts are determined by projecting the residuals onto the derivative of a high-signal-to-noise ratio (SNR) template. Although originally developed for the pRV field, this method is applicable to any spectroscopic time series. The Bouchy et al. (2001) formalism also provides associated uncertainties (see their Equation (13)).
Given the low spectral resolution of SOSS and its limited constraints on radial velocities, we do not attempt to divide the domain into wavelength bins as is done with flux variability. Instead, we consider the integral of Equation (9) from Bouchy et al. (2001) across the 1.00–2.35 m range. The redder portion of the spectrum is excluded due to its markedly different lightcurve behavior.
Figure 8 shows the resulting radial velocity time series, revealing correlated variations at the km s-1 level, characterized by a three-harmonic signal. Per-point uncertainties are approximately 1 km s-1. The signal is in phase with the time derivative of the flux, consistent with the framework described by Aigrain et al. (2012), in which radial velocity (RV) variability for a rotating object (neglecting convective inhibition) scales with the product of the flux and its time derivative—the so-called term in activity-related RV corrections.
@@ IX Doppler Tomography
These results underscore an unexplored opportunity in JWST time-series spectroscopy of brown dwarfs. The high precision of molecular-band flux variability measurements—informing cloud structures at different altitudes—can be complemented by RV-based constraints on longitudinal and latitudinal features. While phase-resolved lightcurves constrain longitudinal structures, RV amplitudes are particularly sensitive to low-latitude features, providing strong priors for brightness map reconstructions. Although the SOSS data for SIMP0136 lacks sufficient SNR to allow subdivision by bandpass, higher-resolution modes such as NIRSpec’s G140H or G235H are of particular interest. For a profile limited by instrumental resolution, RV uncertainty scales with the 3/2 power of resolution at a fixed flux level. Thus, moving from (SOSS) to (G140H/G235H) yields a theoretical sevenfold improvement in SNR, enabling subdivision into several tens of spectral segments while retaining kilometer-per-second-level RV precision.
X Discussion and conclusion
In this work, we presented a multifaceted analysis of time-resolved spectroscopy of the planetary mass object SIMP J01365662+093347. The data were acquired with JWST NIRISS/SOSS and reduced with SOSSISSE. As expected from JWST’s stability, and amply demonstrated with transit work, the data are far beyond what could be obtained through ground-based observations of brown dwarfs. This level of detail opens a new window into brown dwarf atmospheric dynamics.
PCA showed that there are at least three different spectral regions rotating in and out of view. Indeed, not a single atmospheric model could fit the observed spectrum, and a mixture of six Diamondback models is needed to adequately fit the time-averaged spectrum. Projected interpolated Diamondback models on the PCP suggest that disk-integrated physical parameters vary in time. These parameters are temperature, metallicity, and sedimentation efficiency. The fact that we have small spectral variations and change in integrated physical parameters in time ( K, , ), suggests that SIMP 0136 has small-scale cloud patches of different temperature, as predicted by brown dwarf GCM results (Tan & Showman, 2021b). Discrepancy in the reported temperatures of Diamondback models does not affect the magnitude of temperature variability, as it only creates a linear bias. The inhomogeneity in the integrated effective metallicity may arise from different amounts of chemical disequilibrium in each spectral region, e.g., differences in the CO/CH4 ratios among cloud layers.
@@ X Discussion and conclusion
Odd harmonics in the lightcurve of SIMP 0136 and the results from spherical harmonics mapping imply north–south asymmetry, combined with the previously noted inclined rotation. The 2D maps inferred from a single rotation remain degenerate. Nevertheless, lightcurve morphology in combination with Sonora Bobcat contribution functions reveal multidimensional structure of observed variability. We find that the spatial distribution of forsterite clouds below the 1-bar level is anti-correlated with the abundances of water and CO at higher altitudes. This is evidenced by a decrease in disk-integrated emissivity from forsterite layers, accompanied by an increase in emissivity from layers associated with water and CO absorption bands. North–south asymmetry of the brightness map in the – domain could directly arise from the spatial distribution of small-scale forsterite cloud patches. It is not guaranteed that this weather pattern is a permanent feature on SIMP 0136: multiperiod back-to-back observations could show how this morphology evolves.
@@
The fact that the meteorology of SIMP 0136 evolves is confirmed by several ground- and space-based observations at different epochs. In (Artigau et al., 2009), -band (m) photometric observations at Observatoire du Mont-Mégantic in 2008 reveal how the power of Fourier modes varies across nights. Phase-folded lightcurves from September 18 and 19 can be well explained by the first two modes, while phase curves of September 16 and 21 require an odd third-mode harmonic. Further -band observations at Lowell Observatory in 2016 (Croll et al., 2016) show consistent single-peak lightcurves across four rotations on the night of November 10. On the night of November 12, SIMP 0136 shows almost no variability in -band, indicating that meteorological features have longitudinally homogenized. The next night, however, sawtooth-like variations are consistently observed again, but in anticorrelation with the night of November 10, which suggests a slightly denser distribution of clouds in one hemisphere. (Plummer et al., 2024) analyzed CFHT -band observations obtained in 2012: the lightcurve from October 14 is described by three Fourier modes; however, the following night’s observations alternated between Y, J, H, and K bands, limiting the cadence and only allowing two modes to be inferred for each band. A phase shift between H-K and the J-H and H-K color indices was found, which might have indicated different cloud layers. Lastly, we compare our observations with JWST GO #3548 NIRSpec/BOTS time-series spectroscopy analyzed by (McCarthy et al., 2025). The binned lightcurve in the 0.9–2.1 m range is well explained by three Fourier modes, with the third mode having equal power as the second mode, just like in our observations. Yet, 37 hours after our observations, the first mode has evolved and increased its contribution, leading to a different lightcurve morphology between the 2 visits, which is not explained by a simple shift in phase.
@@
We explored a method of extracting and retrieving spectra using spherical harmonics mapping at different wavelengths. This method is valuable in inferring chemical abundances and thermal profiles of potential hot and cold spots of the brightness maps. We performed spectral retrievals on six regions of interest, but the results are not consistent. This is due to the degeneracy of 2D mapping. Observations spanning multiple rotations or Doppler tomography could decrease these degeneracies.
@@
We also use JWST medium resolution spectroscopy for Doppler tomography of SIMP 0136. These measurements could eventually help constrain brightness maps and improve mapping fidelity. The search for a discrepancy between Doppler signals and lightcurve Fourier modes would benefit from higher spectral resolution observations.
Facility: JWST - James Webb Space Telescope (NIRISS).